OpenStudy (anonymous):

Find the value of Q3 for the following data set. 4 6 8 8 8 9 13 15

OpenStudy (anonymous):

x(n+1)/x That's the formula... I'm sorry i don't know this well enough. Send someone like @Hero a message for help! He most probably knows this!

OpenStudy (anonymous):

did u already send it to him?

OpenStudy (anonymous):

No. I kind of thought you would do it...

hero (hero):

I don't even know what Q3 is

hero (hero):

Are you sure you have given all of the information of the problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is a Q3?

OpenStudy (anonymous):

I've heard of it and I'm learning it right now but all i know is the formula

OpenStudy (anonymous):

i have no idea

hero (hero):

I figured it out Q3 is the middle term in the second half of an ordered data set.

hero (hero):

For example, if you have the data set 1 2 3 4 5 6 7 8, The second half of the data set would be 5 6 7 8, and the middle number would be (6 + 7)/2 or 6.5

hero (hero):

In this case, we have the ordered data set 4 6 8 8 8 9 13 15, and the second half is 8 9 13 15. Therefore Q3 = (9 + 13)/2 = 22/2 = 11. So Q3 = 11

hero (hero):

Sorry it took so long. Internet trouble.

OpenStudy (anonymous):

YAY! Now I know it too! Go @Hero!

hero (hero):

But now I'm wondering what Q2 is

OpenStudy (anonymous):

Same... Is it the beginning of the data?

hero (hero):

no that's q1